5x+10=15
6x-12=18
2x+11=21
6x-16=20
4x+12=24
: 5x+10=15
6x-12=18
2x+11=21
6x-16=20
4x+12=24

Answer by nabla(409)

5x+10=15
x=1
6x-12=18
x=5

2x+11=21
x=5
6x-16=20
x=6

4x+12=24
x=3
If you have difficulty arriving at these answers, please include that in your next submission.